The Orion Corp. is evaluating a proposal for a new project. It will cost $50,000 to get the undertaking started. The project will then generate cash inflows of $20,000 in its first year and $16,000 per year in the next five years, after which it will end. Orion uses an interest rate of 19% compounded annually for such evaluations. Calculate the "net present value" (NPV) of the project by treating the initial cost as a cash outflow (a negative) in the present, and adding the present value of the subsequent cash inflows as positives. Round PVF and PVFA values in intermediate calculations to four decimal places. Do not round other intermediate calculations. Round the answer to 2 decimal places. $ What is the implication of a positive NPV? (Words only.) Suppose the inflows were somewhat lower, and the NPV turned out to be negative. What would be the implication of that result? (Words only.)
Answer #1

Subject: Date Cash flow y -_-__ 20.000 16000 16000 16000 16000 16000 - 50,000 NPU= FV (1+0) na no. of period era interest rat

NPV=PVo+PV1+PV2+PV3+PV4+PV5+PV6

=(-50000)+16806.7227+11298.6371+9494.6530+7978.7000+6704.7899+5634.2772

=7917.78

Positive NPV indicates cash inflow is more than cash outflow,means project is profitable while negative NPV indicates that it is not profitable because cash out flow is more than cash inflow.